# Fast Tuning the Index Construction Parameters of Proximity Graphs in Vector Databases

### v1

@@ II Preliminaries
Input : A PG $G$, query $q$, $k$ for $k$-ANN, the entering point $ep$, and a parameter efef for pool size
@@ 4 $u\leftarrow pool[i]$;
8  sort poolpool and keep the $ef$ closest neighbors;
@@ 9 $i\leftarrow$ index of the first unexpanded point in poolpool;
Although various PGs have different graph structures, they share the same $k$-ANNS algorithm, as shown in Algorithm 1. It starts from an entry point $ep$ (line 2), which is specified in advance or randomly selected, and greedily approaches the query $q$. A sorted array poolpool of size $ef\geq k$, which contains the currently found closest neighbors, is maintained and takes $ep$ as its first member (line 2). The main loop (lines 3-8) iteratively extracts the closest unexpanded neighbor $u$ from poolpool (line 4å) and then expands it (lines 5-7), where each $v\in N_{G}(u)$ is treated as a $k$-ANN candidate and verified by an expensive distance computation. Once all the $ef$ nodes in poolpool have been expanded, the search terminates (line 3) and then returns the first $k$ neighbors in poolpool as the $k$-ANN results of $q$ (line 9).
@@ 10     $PN\leftarrow PN\cup\{v\}$;
RNG methods such as NSG [13], DPG [23], NSSG [12] and Vamana [37] assume that they have the full knowledge $D$ in advance. As illustrated in Figure 2a, their construction could be divided into four operations, i.e., Initialization, Search, Prune and Connect. Initialization builds an initial PG $G_{0}$, which could be a KNNG with moderate accuracy as in NSG [13, 23, 12, 35] or a randomly generated KNNG as in Vamana [37]. Then, Search conducts $k$-ANNS (Algorithm 1) for each $u\in D$ with $G_{0}$ in order to obtain refined $k$-ANN results, denoted as $C(u)$. In Prune, each $C(u)$ is pruned to generate $N_{G}(u)$. The most widely used pruning method is the RNG pruning, where $(u,v)\in E$ iff there exists no edge $(u,w)\in E$ such that $\delta(u,w)<\delta(u,v)$ and $\delta(v,w)<\alpha\cdot\delta(u,v)$. $\alpha$ is set as 1 in NSG and HNSW, but is a configurable parameter in Vamana. To be specific, we present the RNG pruning in Algorithm 2. The pruned neighbor set $PN$ for $C(u)$ that is sorted in the ascending order of the distance to $u$, is initialized as $\emptyset$ (line 1). Then, each neighbor $v\in C(u)$ is tried to insert into $PN$ (lines 2-9). $Flag$ is initialized as false (line 3), which indicates whether or not a neighbor in $PN$ dominates $v$. If $\delta(v,w)<\alpha\cdot\delta(u,v)$ (line 5), we say that $w\in PN$ dominates $v$, and set $Flag$ as true (line 6). $v$ will be inserted into $PN$ if all neighbors in $PN$ do not dominate $v$ (lines 7-8). Once $PN$ contains $M$ members, the pruning operation terminates (line 9). As to Connect, it aims to enhance the connectivity of the built PG by adding reversing edges or additional edges between connected components in the current PG.
@@ II-B Problem Statement
Let $N(q)$ denote the set of $k$-ANNS results, and $N^{*}(q)$ the ground truth, $Recall@k={|N(q)\cap N^{*}(q)|}/{k}$. Formally, we give the problem statement as follows. Given an integer $k$ of returned neighbors and a target $Recall@k$ value $t$, our objective is to tune the construction parameters of PG and the $k$-ANNS parameter $ef$ (as in Algorithm 1) to achieve the $k$-ANNS performance of $Recall@k=t$. As aforementioned, we focus on three PGs in this work, i.e., HNSW, Vamana and NSG. (i) HNSW has two construction parameters, i.e., (1) efcefc indicates the parameter $ef$ of $k$-ANNS for each $u\in D$ in Search, and (2) $M$ indicates the out-degree limit in Prune. (ii) Vamana requires four construction parameters, i.e., (1) $L$ indicates $ef$ in $k$-ANNS during Search, (2) $R$ is the number of returned neighbors in Search, (3) $M$ is the out-degree limit in Prune, and (4) $\alpha$ is a parameter in Prune, i.e., $\alpha$ of Algorithm 2. (iii) NSG needs four construction parameters, i.e., (1) $K$ indicates the out-degree of the initial graph in Initialization, (2-4) $L,R,M$ has the same meaning as in Vamana. Notably, our method can be easily extended to other PGs, given their inherent similarity in construction [42].
@@ IV-A Parameter $R$ in RNG Construction
In this part, we analyze the parameter $R$ within RNG construction, as utilized in both NSG and Vamana. This parameter controls the number of neighbors returned in the Search phase for subsequent Prune. To elaborate, consider $pool_{L}=\{v_{1},v_{2},\cdots,v_{L}\}$ as the data point set obtained from $k$-ANNS (Algorithm 1) when $ef=L$, with $q$ is a data point $u\in D$. Assuming $\delta(v_{1},u)\leq\delta(v_{2},u)\leq\cdots\leq\delta(v_{L},u)$, $R$ is employed to establish $C(u)=\{v_{1},v_{2},\cdots,v_{R}\}$, where $R\leq L$. Let $PN(R)$ denote the set $PN$ derived from the Prune process (Algorithm 2) when $|C(u)|=R$ for a node $u$ and $C(u)$ is conducted with the same value of $L$ and same PG. The subsequent theorem demonstrates that the set $PN(R)$ is a subset of $PN(R^{\prime})$ when $R\leq R^{\prime}$.
@@
If $|PN(R)|\geq M$, the algorithm returns $PN$ upon termination at line 9. Hence, elements $v\in C(u)$ with greater distances to $u$ are not enumerated. This results in $PN(R)=PN(R^{\prime})$ for $R^{\prime}\geq R$. Otherwise, if we let $PN(R)={v_{p_{1}},v_{p_{2}},\cdots,v_{p_{c}}}$, where $c<M$ and $p_{1}<p_{2}<\cdots<p_{c}$, for $i<p_{c}$ such that $v_{i}\not\in PN(R)$, it follows that $v_{i}\not\in PN(R^{\prime})$ as well. This is because both $PN(R)$ and $PN(R^{\prime})$ operate on the same elements in the same order for pruning nodes, if a node $v_{j}\in PN(R)$ satisfies $\alpha\cdot\delta(v_{i},v_{j})<\delta(v_{i},u)$, it will also exclude $v_{i}$ from $PN(R^{\prime})$. ∎
Hence, from the above theorem, it is evident that increasing the value of $R$ initially results in a sufficient neighbor list, i.e., the $|PN|$ increases and finally reaches $M$, and then will not affect $PN$. Additionally, since the efficiency of Algorithm 2 remains unaffected by the value of $R$, courtesy of the early break at line 9, the traversal of more candidate nodes in $C(u)$ is avoided when the neighbor size is $M$. Therefore, setting the value of $R$ as $L$ during index construction is a direct approach, rendering fine-tuning unnecessary, as opting for $R=L$ proves to be the optimal choice.
@@ IV-C Parameter Estimation Analysis
If $|PN(M)|<M$, all elements $v\in C(u)$ have been traversed. Thus, the value of $PN$ is independent of $M$, leading to $PN(M)=PN(M^{\prime})$. If $|PN(M)|\geq M$, the algorithm returns $PN$ at line 9 upon termination. Hence, elements $v\in C(u)$ with greater distances to $u$ are not enumerated for $M$, but they may be included with a larger value of $M$ and could be contained in $PN(M^{\prime})$. This implies $PN(M)\subseteq PN(M^{\prime})$ for $M^{\prime}\geq M$. ∎
Next, we analyze the size $|C(u)|$ of the candidate neighbor set for Prune, i.e., efcefc in HNSW and $L$ in NSG and Vamana, as well as $\alpha$. We analyze these parameters in an experimental study to demonstrate their effects on the overlaps between the graph structures. Here, we take Vamana as an example. For two PGs $G_{1}$ and $G_{2}$ with different parameters, the overlap of two neighbor lists on $u$, i.e., $N_{G_{1}}(u)$ and $N_{G_{2}}(u)$, is defined as $NLO(N_{G_{1}}(u),N_{G_{2}}(u))=|N_{G_{1}}(u)\cap N_{G_{2}}(u)|/|N_{G_{1}}(u)|$. Hence, the overlap between $G_{1}$ and $G_{2}$ is defined as $NLO(G_{1},G_{2})=1/|D|\cdot\sum_{u\in D}NLO(N_{G_{1}}(u),N_{G_{2}}(u))$. We show the results in Figure 5, where we fix $M=50$. In Figure 5a, we fix $\alpha=1.2$ and present the $NLO$ value between two graphs with different $L$ values. Obviously, the closer $L$ values are, the larger the $NLO$ is, i.e., the more similar their graph structures are. In Figure 5b, we fix $L=100$ and vary $\alpha$. Similarly, the closer $\alpha$ values lead to similar graph structures built accordingly. Besides, similar phenomena could be observed in other PGs and other datasets, which are omitted due to the space limit.
@@ IV-D Efficient Search Operations for Multiple PGs
To illustrate this phenomenon, we take HNSW as an example and conduct experiments in the real-world datasets, whose results are shown in Table II. Here, we build three HNSW graphs with three construction parameters and count the repeated distance computations in the construction of those HNSW graphs. Let $dist_{A}$, $dist_{B}$, and $dist_{C}$ be the set of vector pairs for distance computations under the parameter settings $A$, $B$, and $C$, respectively. We use #dist\#dist to denote the total distance computations, and $ratio_{rp}$ to denote the ratio of common distance computations among all three settings, i.e., $ratio_{rp}=|dist_{A}\cap dist_{B}\cap dist_{C}|/(|dist_{A}|+|dist_{B}|+|dist_{C}|)$. Let $dist_{A}^{s}$, $dist_{B}^{s}$, and $dist_{C}^{s}$ be the set of vector pairs for distance computations caused by Search operations. $ratio_{rp}^{s}=|dist_{A}^{s}\cap dist_{B}^{s}\cap dist_{C}^{s}|/(|dist_{A}^{s}|+|dist_{B}^{s}|+|dist_{C}^{s}|)$. The results show that more than half of the distance computations are repeated under the different parameter settings. Regarding Search operations, they share a substantial portion (i.e., at least 60%) of redundant distance computations, excluding those that could be prevented by utilizing a bitmap visitedvisited in Algorithm 1, where each element in visitvisit indicates whether or not it has been verified in this $k$-ANNS operation.
Input : A PG $G$, query $q$, $k$ for $k$-ANN, the entering point $ep$, a parameter efef for pool size, and $V_{\delta}$ for record distance computations
@@ 4 $u\leftarrow pool[i]$;
11  sort poolpool and keep the $ef$ closest neighbors;
@@ IV-F Efficient Construction of Multiple PGs
Efficient Construction of Multiple HNSW Graphs. We present the details in Algorithm 5. Notably, $ef$ and $M$ are two key construction parameters of HNSW. First, we initialize each HNSW graph (lines 1-3), where $m_{L}$ indicates the highest layer of the current HNSW graphs. Then, we insert each $u\in D$ into the graphs iteratively (lines 4-19). In the loop, we first determine the highest layer $l$ for $u$ (line 5), which is used for all $m$ graphs, and then update $m_{L}$ if necessary (line 6). We initialize $V_{\delta}$ (line 7) and then insert $u$ into the $m$ graphs one by one (lines 8-19). For the layers from $m_{L}$ to $l+1$, we find the 1-ANN of $u$ in each layer, and use it as the entry point of the next layer (line 11). Here, $G_{i}^{j}$ represents the $j$-th layer of the $i$-th graph $G_{i}$. For the remaining layers from $l$ to 0, we find their $efc_{j}$-ANNs using our efficient mKANNS method (line 14), followed by our mPrune method (line 15). Then, we add reverse edges from the pruned neighbors to $u$ (lines 17) and ensure that the node’s out-degree limit is met (lines 18-19). Note that this method could be easily extended to other NSWGs such as NSW [30], where we fix $m_{L}$ as 1, employ the closeness-first pruning method, and remove the out-degree limit. Note that $V_{\delta}$ will be deleted after inserting $u$ into all $m$ graphs, since storing $n$ $V_{\delta}$ arrays requires $O(n^{2})$ space.

### v2

@@ II Preliminaries
Input : A PG $G$, query $q$, $k$ for $k$-ANN, the entering point epep, and a parameter efef for pool size
@@ 4 $u\leftarrow pool[i]$;
8  sort poolpool and keep the efef closest neighbors;
@@ 9 $i\leftarrow$ index of the first unexpanded point in poolpool;
Although various PGs have different graph structures, they share the same $k$-ANNS algorithm, as shown in Algorithm 1. It starts from an entry point epep (line 2), which is specified in advance or randomly selected, and greedily approaches the query $q$. A sorted array poolpool of size $ef\geq k$, which contains the currently found closest neighbors, is maintained and takes epep as its first member (line 2). The main loop (lines 3-8) iteratively extracts the closest unexpanded neighbor $u$ from poolpool (line 4å) and then expands it (lines 5-7), where each $v\in N_{G}(u)$ is treated as a $k$-ANN candidate and verified by an expensive distance computation. Once all the efef nodes in poolpool have been expanded, the search terminates (line 3) and then returns the first $k$ neighbors in poolpool as the $k$-ANN results of $q$ (line 9).
@@ 10     $PN\leftarrow PN\cup\{v\}$;
RNG methods such as NSG [13], DPG [23], NSSG [12] and Vamana [37] assume that they have the full knowledge $D$ in advance. As illustrated in Figure 2a, their construction could be divided into four operations, i.e., Initialization, Search, Prune and Connect. Initialization builds an initial PG $G_{0}$, which could be a KNNG with moderate accuracy as in NSG [13, 23, 12, 35] or a randomly generated KNNG as in Vamana [37]. Then, Search conducts $k$-ANNS (Algorithm 1) for each $u\in D$ with $G_{0}$ in order to obtain refined $k$-ANN results, denoted as $C(u)$. In Prune, each $C(u)$ is pruned to generate $N_{G}(u)$. The most widely used pruning method is the RNG pruning, where $(u,v)\in E$ iff there exists no edge $(u,w)\in E$ such that $\delta(u,w)<\delta(u,v)$ and $\delta(v,w)<\alpha\cdot\delta(u,v)$. $\alpha$ is set as 1 in NSG and HNSW, but is a configurable parameter in Vamana. To be specific, we present the RNG pruning in Algorithm 2. The pruned neighbor set PNPN for $C(u)$ that is sorted in the ascending order of the distance to $u$, is initialized as $\emptyset$ (line 1). Then, each neighbor $v\in C(u)$ is tried to insert into PNPN (lines 2-9). FlagFlag is initialized as false (line 3), which indicates whether or not a neighbor in PNPN dominates $v$. If $\delta(v,w)<\alpha\cdot\delta(u,v)$ (line 5), we say that $w\in PN$ dominates $v$, and set FlagFlag as true (line 6). $v$ will be inserted into PNPN if all neighbors in PNPN do not dominate $v$ (lines 7-8). Once PNPN contains $M$ members, the pruning operation terminates (line 9). As to Connect, it aims to enhance the connectivity of the built PG by adding reversing edges or additional edges between connected components in the current PG.
@@ II-B Problem Statement
Let $N(q)$ denote the set of $k$-ANNS results, and $N^{*}(q)$ the ground truth, $Recall@k={|N(q)\cap N^{*}(q)|}/{k}$. Formally, we give the problem statement as follows. Given an integer $k$ of returned neighbors and a target $Recall@k$ value $t$, our objective is to tune the construction parameters of PG and the $k$-ANNS parameter efef (as in Algorithm 1) to achieve the $k$-ANNS performance of $Recall@k=t$. As aforementioned, we focus on three PGs in this work, i.e., HNSW, Vamana and NSG. (i) HNSW has two construction parameters, i.e., (1) efcefc indicates the parameter efef of $k$-ANNS for each $u\in D$ in Search, and (2) $M$ indicates the out-degree limit in Prune. (ii) Vamana requires four construction parameters, i.e., (1) $L$ indicates efef in $k$-ANNS during Search, (2) $R$ is the number of returned neighbors in Search, (3) $M$ is the out-degree limit in Prune, and (4) $\alpha$ is a parameter in Prune, i.e., $\alpha$ of Algorithm 2. (iii) NSG needs four construction parameters, i.e., (1) $K$ indicates the out-degree of the initial graph in Initialization, (2-4) $L,R,M$ has the same meaning as in Vamana. Notably, our method can be easily extended to other PGs, given their inherent similarity in construction [42].
@@ IV-A Parameter $R$ in RNG Construction
In this part, we analyze the parameter $R$ within RNG construction, as utilized in both NSG and Vamana. This parameter controls the number of neighbors returned in the Search phase for subsequent Prune. To elaborate, consider $pool_{L}=\{v_{1},v_{2},\cdots,v_{L}\}$ as the data point set obtained from $k$-ANNS (Algorithm 1) when $ef=L$, with $q$ is a data point $u\in D$. Assuming $\delta(v_{1},u)\leq\delta(v_{2},u)\leq\cdots\leq\delta(v_{L},u)$, $R$ is employed to establish $C(u)=\{v_{1},v_{2},\cdots,v_{R}\}$, where $R\leq L$. Let $PN(R)$ denote the set PNPN derived from the Prune process (Algorithm 2) when $|C(u)|=R$ for a node $u$ and $C(u)$ is conducted with the same value of $L$ and same PG. The subsequent theorem demonstrates that the set $PN(R)$ is a subset of $PN(R^{\prime})$ when $R\leq R^{\prime}$.
@@
If $|PN(R)|\geq M$, the algorithm returns PNPN upon termination at line 9. Hence, elements $v\in C(u)$ with greater distances to $u$ are not enumerated. This results in $PN(R)=PN(R^{\prime})$ for $R^{\prime}\geq R$. Otherwise, if we let $PN(R)={v_{p_{1}},v_{p_{2}},\cdots,v_{p_{c}}}$, where $c<M$ and $p_{1}<p_{2}<\cdots<p_{c}$, for $i<p_{c}$ such that $v_{i}\not\in PN(R)$, it follows that $v_{i}\not\in PN(R^{\prime})$ as well. This is because both $PN(R)$ and $PN(R^{\prime})$ operate on the same elements in the same order for pruning nodes, if a node $v_{j}\in PN(R)$ satisfies $\alpha\cdot\delta(v_{i},v_{j})<\delta(v_{i},u)$, it will also exclude $v_{i}$ from $PN(R^{\prime})$. ∎
Hence, from the above theorem, it is evident that increasing the value of $R$ initially results in a sufficient neighbor list, i.e., the $|PN|$ increases and finally reaches $M$, and then will not affect PNPN. Additionally, since the efficiency of Algorithm 2 remains unaffected by the value of $R$, courtesy of the early break at line 9, the traversal of more candidate nodes in $C(u)$ is avoided when the neighbor size is $M$. Therefore, setting the value of $R$ as $L$ during index construction is a direct approach, rendering fine-tuning unnecessary, as opting for $R=L$ proves to be the optimal choice.
@@ IV-C Parameter Estimation Analysis
If $|PN(M)|<M$, all elements $v\in C(u)$ have been traversed. Thus, the value of PNPN is independent of $M$, leading to $PN(M)=PN(M^{\prime})$. If $|PN(M)|\geq M$, the algorithm returns PNPN at line 9 upon termination. Hence, elements $v\in C(u)$ with greater distances to $u$ are not enumerated for $M$, but they may be included with a larger value of $M$ and could be contained in $PN(M^{\prime})$. This implies $PN(M)\subseteq PN(M^{\prime})$ for $M^{\prime}\geq M$. ∎
Next, we analyze the size $|C(u)|$ of the candidate neighbor set for Prune, i.e., efcefc in HNSW and $L$ in NSG and Vamana, as well as $\alpha$. We analyze these parameters in an experimental study to demonstrate their effects on the overlaps between the graph structures. Here, we take Vamana as an example. For two PGs $G_{1}$ and $G_{2}$ with different parameters, the overlap of two neighbor lists on $u$, i.e., $N_{G_{1}}(u)$ and $N_{G_{2}}(u)$, is defined as $NLO(N_{G_{1}}(u),N_{G_{2}}(u))=|N_{G_{1}}(u)\cap N_{G_{2}}(u)|/|N_{G_{1}}(u)|$. Hence, the overlap between $G_{1}$ and $G_{2}$ is defined as $NLO(G_{1},G_{2})=1/|D|\cdot\sum_{u\in D}NLO(N_{G_{1}}(u),N_{G_{2}}(u))$. We show the results in Figure 5, where we fix $M=50$. In Figure 5a, we fix $\alpha=1.2$ and present the NLONLO value between two graphs with different $L$ values. Obviously, the closer $L$ values are, the larger the NLONLO is, i.e., the more similar their graph structures are. In Figure 5b, we fix $L=100$ and vary $\alpha$. Similarly, the closer $\alpha$ values lead to similar graph structures built accordingly. Besides, similar phenomena could be observed in other PGs and other datasets, which are omitted due to the space limit.
@@ IV-D Efficient Search Operations for Multiple PGs
To illustrate this phenomenon, we take HNSW as an example and conduct experiments in the real-world datasets, whose results are shown in Table II. Here, we build three HNSW graphs with three construction parameters and count the repeated distance computations in the construction of those HNSW graphs. Let $dist_{A}$, $dist_{B}$, and $dist_{C}$ be the set of vector pairs for distance computations under the parameter settings AA, $B$, and $C$, respectively. We use #dist\#dist to denote the total distance computations, and $ratio_{rp}$ to denote the ratio of common distance computations among all three settings, i.e., $ratio_{rp}=|dist_{A}\cap dist_{B}\cap dist_{C}|/(|dist_{A}|+|dist_{B}|+|dist_{C}|)$. Let $dist_{A}^{s}$, $dist_{B}^{s}$, and $dist_{C}^{s}$ be the set of vector pairs for distance computations caused by Search operations. $ratio_{rp}^{s}=|dist_{A}^{s}\cap dist_{B}^{s}\cap dist_{C}^{s}|/(|dist_{A}^{s}|+|dist_{B}^{s}|+|dist_{C}^{s}|)$. The results show that more than half of the distance computations are repeated under the different parameter settings. Regarding Search operations, they share a substantial portion (i.e., at least 60%) of redundant distance computations, excluding those that could be prevented by utilizing a bitmap visitedvisited in Algorithm 1, where each element in visitvisit indicates whether or not it has been verified in this $k$-ANNS operation.
Input : A PG $G$, query $q$, $k$ for $k$-ANN, the entering point epep, a parameter efef for pool size, and $V_{\delta}$ for record distance computations
@@ 4 $u\leftarrow pool[i]$;
11  sort poolpool and keep the efef closest neighbors;
@@ IV-F Efficient Construction of Multiple PGs
Efficient Construction of Multiple HNSW Graphs. We present the details in Algorithm 5. Notably, efef and $M$ are two key construction parameters of HNSW. First, we initialize each HNSW graph (lines 1-3), where $m_{L}$ indicates the highest layer of the current HNSW graphs. Then, we insert each $u\in D$ into the graphs iteratively (lines 4-19). In the loop, we first determine the highest layer $l$ for $u$ (line 5), which is used for all $m$ graphs, and then update $m_{L}$ if necessary (line 6). We initialize $V_{\delta}$ (line 7) and then insert $u$ into the $m$ graphs one by one (lines 8-19). For the layers from $m_{L}$ to $l+1$, we find the 1-ANN of $u$ in each layer, and use it as the entry point of the next layer (line 11). Here, $G_{i}^{j}$ represents the $j$-th layer of the $i$-th graph $G_{i}$. For the remaining layers from $l$ to 0, we find their $efc_{j}$-ANNs using our efficient mKANNS method (line 14), followed by our mPrune method (line 15). Then, we add reverse edges from the pruned neighbors to $u$ (lines 17) and ensure that the node’s out-degree limit is met (lines 18-19). Note that this method could be easily extended to other NSWGs such as NSW [30], where we fix $m_{L}$ as 1, employ the closeness-first pruning method, and remove the out-degree limit. Note that $V_{\delta}$ will be deleted after inserting $u$ into all $m$ graphs, since storing $n$ $V_{\delta}$ arrays requires $O(n^{2})$ space.